OpenStudy (rockstar0765):

The probability that a given urine sample analyzed in a pathology lab contains protein is 0.73 (event A). The probability that the urine sample contains the protein albumin is 0.58 (event B). The probability that a sample contains protein, given that albumin is detected, is 1. Which statement is true?

OpenStudy (rockstar0765):

Events A and B are dependent because P(A|B) = P(A). Events A and B are dependent because P(A| B) P(B). Events A and B are dependent because P(A|B) = P(A) + P(B). Events A and B are dependent because P(A| B) P(A). Events A and B are dependent because P(A|B) = P(A).

OpenStudy (amistre64):

how do we define dependence? or rather independence?

OpenStudy (amistre64):

if P(A|B) = P(A|C) = P(A|D) = ... = P(A|N) then the value of P(A) does not depend on the even in which it is occuring.

OpenStudy (amistre64):

**on the event .....

OpenStudy (amistre64):

if there is not an equality, then the probability of A is NOT the same for all given events, and therefore it depends on which event A takes place in. dependence is therefore attributed to an inequality of the probability of A, among the events in which A occurs.

OpenStudy (kobeni-chan):

Are the blank spaces in the second and fourth answers, that n looking symbol and a (not equal) sign?